### v1

# Jensen’s Functional Equation on Involution-Generated Groups: An ($\mathrm{SR}_{2}$) Criterion and Applications

Abstract.
We study the Jensen functional equations on a group $G$ with values in an abelian group $H$:
(J1) $f(xy)+f(xy^{-1})$ $=2f(x)\qquad(\forall\,x,y\in G),$
(J2) $f(xy)+f(x^{-1}y)$ $=2f(y)\qquad(\forall\,x,y\in G),$
with the normalization $f(e)=0$. Building on techniques for the symmetric groups $S_{n}$, we isolate a structural criterion on GG—phrased purely in terms of involutions and square roots—under which every solution to (J1) must also satisfy (J2) and is automatically a group homomorphism. Our new criterion, denoted $(\mathrm{SR}_{2})$, implies that $S_{1}(G,H)=S_{1,2}(G,H)=\mathrm{Hom}(G,H)$, applies to many reflection–generated groups and, in particular, recovers the full solution on $S_{n}$. Furthermore, we give a transparent description of the solution space in terms of the abelianization $G/[G,G]$, and we treat dihedral groups $D_{m}$ in detail, separating the cases $m$ odd vs. even. The approach is independent of division by $2$ in $H$ and complements the classical complex-valued theory that reduces (J1) to functions on $G/[G,[G,G]]$.
Key words and phrases:
Jensen’s functional equation, groups generated by involutions, group homomorphisms, symmetric groups, dihedral groups.
2020 Mathematics Subject Classification:
Primary 39B52; Secondary 20F55.
1. Introduction
On the real line, Jensen’s equation captures convexity. On noncommutative groups, two canonical Jensen–type equations emerge:
(J1) $f(xy)+f(xy^{-1})$ $=2f(x)$
(J2) $f(xy)+f(x^{-1}y)$ $=2f(y).$
A systematic study on groups was initiated by C. T. Ng [3, 4, 5], who developed reduction formulas and solved Pexiderized variants on important classes (free groups, linear groups, semidirect products). Remarkably, in [6], Stetkær showed that for complex-valued solutions of (J1), every solution factors through the $2$-step derived quotient
$G\longrightarrow G/[G,[G,G]],$
and provided explicit solution formulas; moreover, the odd-solution quotient by homomorphisms is canonically isomorphic to $\mathrm{Hom}([G,G]/[G,[G,G]],\mathbb{C})$.
In parallel, Jensen-type equations on (possibly noncommutative) semigroups with endomorphisms have recently been solved under 2-torsion-free hypotheses on the codomain, dropping involutivity of endomorphisms and expressing solutions via additive maps [2]; see also [1] relating Jensen-type and quadratic equations with endomorphisms. These developments largely rely either on divisibility by $2$ in the codomain or on endomorphism structure in the domain.
Motivation and contribution
The present work is motivated by the specific case of the symmetric group $S_{n}$. The result that all Jensen solutions on $S_{n}$ are homomorphisms was previously asserted by C.T. Ng in his foundational work [4, 5], though a detailed proof was not published. Addressing this, Trinh and Hieu [7] later provided the first direct and elementary proof of this important fact. The key observation in [7] is that the product of any two transpositions in $S_{n}$ is always a square. This property leads to two interesting consequences for any Jensen solution $f$: first that $2f\equiv 0$, and second that $f$ is invariant under reordering of factors, from which additivity ($f(xy)=f(x)+f(y)$) follows.
In this paper, we propose a new structural criterion, which we call ”square-root for products of involutions” $(\mathrm{SR}_{2})$. We prove that this criterion has a beautiful consequence: for any group satisfying $(\mathrm{SR}_{2})$, every solution $f$ to the first Jensen equation (J1) (with $f(e)=0$) must also satisfy the second (J2) and be a group homomorphism. This result is highly general, as our proof avoids the common technical assumptions on the codomain group $H$ (such as divisibility by 2). The criterion’s power is demonstrated by its broad applicability, providing a unified explanation for the symmetric group case and extending to a class of ”reflection-generated” groups—groups of symmetries constructed from geometric flips. This class includes well-known examples such as certain dihedral groups and their important generalizations, the Coxeter groups.
Building on the arrived results, we then identify the solution spaces $S_{1}(G,H)$ and $S_{1,2}(G,H)$ as being identical to the group $\mathrm{Hom}(G/[G,G],H)$, providing an immediate combinatorial description whenever $G/[G,G]$ is a 2-group. Finally, we give a sharp dichotomy for the dihedral groups $D_{m}$, showing that the mechanism works exactly when $m$ is odd.
Relation to prior work.
Unlike the complex-valued theory in [6], our approach does not divide by $2$ in $H$; unlike recent semigroup results [2, 1], our criterion $(\mathrm{SR}_{2})$ is a purely group-structural hypothesis on $G$ (no endomorphisms required) and targets the full equality $S_{1}(G,H)=S_{1,2}(G,H)=\mathrm{Hom}(G,H)$, showing this follows from (J1) alone.
Organization of the paper
The paper is organized as follows. In Section 2, we introduce our main structural condition, the $(\mathrm{SR}_{2})$ property, and develop the necessary technical lemmas (Lemmas 2.32.7). We then prove the main result (Theorem 2.8), which states that for any group satisfying $(\mathrm{SR}_{2})$, every solution to (J1) is a group homomorphism. In Section 3, we apply this theorem to key examples: we recover the known results for the symmetric groups $S_{n}$ (Theorem 3.1) and provide a detailed analysis of the dihedral groups $D_{m}$, establishing a sharp dichotomy between the $m$ odd and $m$ even cases (Theorem 3.2). This section includes an explicit counterexample (Example 3.4) for the even case. Finally, in Section 4, we compare our approach to the classical complex-valued theory and recent results on semigroups.
Notation
Throughout this paper, the group $G$ is written multiplicatively, while the abelian group $H$ is written additively. We also normalize our function by assuming that $f(e)=0.$
For a group $G$, write $S_{1}(G,H)$ (resp. $S_{2}(G,H)$) for the solution sets of (J1) (resp. (J2)). Then, $S_{1,2}(G,H):=S_{1}(G,H)\cap S_{2}(G,H).$
We write
$G_{\mathrm{ab}}:=G/[G,G]$
for the abelianization of $G$, and denote by $\pi:G\to G_{\mathrm{ab}}$ the canonical projection (any homomorphism from $G$ to an abelian group factors uniquely through $\pi$).
2. A structural square-root criterion
Definition 2.1 (The $(\mathrm{SR}_{2})$ property).
Let $G$ be a group and $\mathcal{I}\subseteq G$ a set of involutions ($i^{2}=e$). We say that $G$ satisfies $\mathrm{SR}_{2}(\mathcal{I})$ if:
  1. (1)
    $G=\langle\mathcal{I}\rangle$ (i.e. $G$ is generated by involutions from $\mathcal{I}$);
  2. (2)
    for every $a,b\in\mathcal{I}$ there exists $t\in G$ with $t^{2}=ab$.
We write simply $(\mathrm{SR}_{2})$ when $\mathcal{I}$ is understood.
Remark 2.2.
For $G=S_{n}$ with $\mathcal{I}$ the set of transpositions, one checks directly that the product of any two transpositions is a square in the subgroup they generate; this is the combinatorial heart of the proof on $S_{n}$: see [7].
Lemma 2.3 (Basic identities without dividing by $2$).
Let $G$ be a group and $H$ an abelian group. Suppose $f:G\to H$ satisfies the two Jensen identity (J1) with $f(e)=0$. Then for all $x,y,z\in G$:
  1. (I)
    Oddness and square rule:
    $f(x^{-1})=-f(x),\qquad f(x^{2})=2f(x).$
  2. (II)
    Three–variable switching formulas (explicit, no $1/2$):
    (2.1) $f(xyz)$ $=2f(x)-f\!\big(xz^{-1}y^{-1}\big),$
    (2.2) $f(xzy)$ $=2f(x)-f\!\big(xy^{-1}z^{-1}\big).$
    Consequently,
    (2.3) $f(xyz)-f(xzy)=f\!\big(xy^{-1}z^{-1}\big)-f\!\big(xz^{-1}y^{-1}\big).$
Proof.
Setting $x=e$ in (J1) gives $f(y)+f(y^{-1})=2f(e)=0$, i.e. $f(y^{-1})=-f(y)$. With $y=x$ in (J1) we get $f(x^{2})+f(e)=2f(x)$, hence $f(x^{2})=2f(x)$ since $f(e)=0$. This proves (I).
For (II), apply (J1) with $(x,\,y)\mapsto\big(x,\,yz\big)$ and with $(x,\,y)\mapsto\big(x,\,zy\big)$ to obtain (2.1)–(2.2). Subtracting (2.2) from (2.1) yields (2.3). ∎
Remark 2.4.
Formulas (2.1) and (2.2) are the precise ”three–variable manipulations” we use later to permute adjacent factors at the level of ff–values, and they never invoke division by 22.
We now present the main results of this work along with their proofs.
Theorem 2.5 (Two involutions: torsion bounds and square roots).
Let $G$ be a group, $H$ an abelian group, and $f:G\to H$ satisfy (J1) with $f(e)=0$. Let $a,b\in G$ be involutions ($a^{2}=b^{2}=e$). Then:
  1. (I)
    $f(a^{-1})=-f(a)$ and $2f(a)=0$; likewise $f(b^{-1})=-f(b)$ and $2f(b)=0$.
  2. (II)
    $2f(ab)=0$.
  3. (III)
    If, in addition, (SR2) holds for $a,b$ (so there is $t\in G$ with $t^{2}=ab$), then
    $f(ab)=f(t^{2})=2f(t)\quad\text{and hence}\quad 4f(t)=0.$
    In particular, if one further knows that $f(ab)=0$ (e.g. from a later lemma establishing pair–vanishing, or when $H$ is $2$-torsion-free), then $2f(t)=0$.
Proof.
(I) Apply Lemma 2.3(I): from (J1) with $x=e$ we get $f(y^{-1})=-f(y)$, and with $y=x$ we obtain $f(x^{2})=2f(x)$. For an involution $a=a^{-1}$ this gives $2f(a)=f(a^{2})=f(e)=0$; similarly for $b$.
(II) Using (J1) at $(x,y)=(a,b)$ and the fact $b^{-1}=b$,
$f(ab)+f(ab^{-1})=2f(a)\quad\Rightarrow\quad 2f(ab)=2f(a)=0$
by part (I).
(III) If (SR2) provides $t$ with $t^{2}=ab$, then Lemma 2.3(I) yields
$f(ab)=f(t^{2})=2f(t).$
Combining with part (II) gives $4f(t)=0$. If moreover $f(ab)=0$ (for instance because a later ”pair–vanishing” lemma shows $f$ vanishes on any product of two involutions, or because $H$ is $2$-torsion-free so from $2f(ab)=0$ we get $f(ab)=0$), then necessarily $2f(t)=0$. ∎
Lemma 2.6.
Let $G$ be a group, $H$ an abelian group, and $f:G\to H$ satisfy (J1) with $f(e)=0$. If $g\in G$ can be written as a product of involutions, then
2f(g)=0.2\,f(g)=0.
In particular, if $G$ is generated by involutions, then $2f\equiv 0$ on $G$.
Proof.
Write $g=i_{1}i_{2}\cdots i_{r}$ with each $i_{k}^{2}=e$. We prove by induction on $r$ that $2f(i_{1}\cdots i_{r})=0$. For $r=0$ (i.e. $g=e$), $2f(g)=0$ since $f(e)=0$. Assume the claim holds for <r<r and set $X=i_{1}\cdots i_{r-1}$ and $a=i_{r}$ (so $a^{2}=e$). Applying (J1) with $(x,y)=(X,a)$ gives
$f(Xa)+f(Xa^{-1})=2f(X).$
Because $a=a^{-1}$, the left-hand side is $2f(Xa)$. Hence
$2f(i_{1}\cdots i_{r})=2f(Xa)=2f(X)=0$
by the induction hypothesis. This completes the induction. ∎
Theorem 2.7 (Reordering invariance under $2f\equiv 0$).
Let $G$ be a group generated by involutions and $H$ an abelian group. Assume $f:G\to H$ satisfies (J1) with $f(e)=0$. If $2f\equiv 0$ on $G,$ then:
  1. (I)
    For all $x,y,z\in G$, one has $f(xyz)=f(xzy)$.
  2. (II)
    Consequently, for any word $g=i_{1}\cdots i_{r}$ in involutions, swapping any adjacent pair $i_{k},i_{k+1}$ does not change the value $f(g)$.
Proof.
(I) By Lemma 2.3(II), we have the explicit switch identities:
$f(xyz)=2f(x)-f\!\big(xz^{-1}y^{-1}\big),\qquad f(xzy)=2f(x)-f\!\big(xy^{-1}z^{-1}\big).$
Subtracting the second identity from the first yields the consequence:
$f(xyz)-f(xzy)=f(xy^{-1}z^{-1})-f(xz^{-1}y^{-1}).$
Applying (J1) to $(xy^{-1},z)$ and $(xz^{-1},y)$ and using the hypothesis $2f\equiv 0$ shows that $f(xy^{-1}z^{-1})=-f(xy^{-1}z)$ and $f(xz^{-1}y^{-1})=-f(xz^{-1}y)$. Hence the difference becomes
$f(xyz)-f(xzy)=-f(xy^{-1}z)+f(xz^{-1}y).$
The switching identity (2.1), $f(uvw)=2f(u)-f(uw^{-1}v^{-1})$, combined with $2f\equiv 0$ yields the rule $f(uvw)=-f(uw^{-1}v^{-1})$. Applying this to the term $f(xy^{-1}z)$ gives
$f(xy^{-1}z)=-f(xz^{-1}(y^{-1})^{-1})=-f(xz^{-1}y).$
Therefore, the expression for the difference simplifies to
$f(xz^{-1}y)+f(xz^{-1}y)=2f(xz^{-1}y),$
which is 0 by Lemma 2.6. Thus $f(xyz)=f(xzy)$, as claimed.
(II) The statement for adjacent swaps in a word follows by taking $x$ to be the prefix before the pair, $y=i_{k}$, $z=i_{k+1}$. ∎
Theorem 2.8.
Let $G$ be a group generated by involutions that satisfies the (SR2) condition, and let $H$ be an arbitrary abelian group. If $f:G\to H$ satisfies (J1) and $f(e)=0$, then $f$ is a group homomorphism and also satisfies (J2).
Consequently, under these conditions on $G$, the solution spaces are identical:
$S_{1}(G,H)=S_{1,2}(G,H)=\mathrm{Hom}(G,H)\cong\mathrm{Hom}(G_{\mathrm{ab}},H).$
Proof.
Fix a generating set $\mathcal{I}$ of involutions of $G$. For $g\in G$ define the involution-word length
$\ell(g):=\min\{\,r\geq 0:\ g=i_{1}\cdots i_{r}\text{ with }i_{k}\in\mathcal{I}\,\}.$
(Existence follows since $G$ is generated by involutions.)
We will prove additivity $f(xy)=f(x)+f(y)$ by induction on $q:=\ell(y)$, keeping $x$ fixed.
Case $q=0$. Trivial since $y=e$ and $f(e)=0$.
Case $q=1$. Let $y=j$ be an involution and define c(j;x):=f(xj)f(x)c(j;x):=f(xj)-f(x).
Step A: 22–torsion. From (J1) with $(x,y)=(x,j)$ and $j^{-1}=j$ we have $2f(xj)=2f(x)$, hence
(2.4) $2\,c(j;x)=0\qquad(\forall\,x\in G).$
Step B: reordering and absorption tools. By Theorem 2.7, $f(Zt^{2}j)=f(Zjt^{2})$ for all $Z,t$. Applying (J1) to $(x,y)=(Zt,t)$ and using Lemma 2.6 ($2f\equiv 0$ on words of involutions) yields the absorption identity
(2.5) $f(Zt^{2})=-\,f(Z)\quad\forall\,Z,\,t\in G.$
Step C: two invariances for $c(j;\cdot)$.
  • (i)
    Right–multiplying by $j$ preserves $c$. Indeed,
    c(j;xj)=f(xjj)f(xj)=f(x)f(xj)=c(j;x).c(j;xj)=f(xjj)-f(xj)=f(x)-f(xj)=-\,c(j;x).
    By (2.4) we have c(j;x)=c(j;x)-c(j;x)=c(j;x), hence $c(j;xj)=c(j;x)$.
  • (ii)
    Right–multiplying by a square $t^{2}$ preserves $c$. Using reordering and equations (2.4), (2.5), we get
    $c(j;xt^{2})$ $=f(xt^{2}j)-f(xt^{2})=f(xjt^{2})-f(xt^{2})$
    $=-\,f(xj)-\big(-\,f(x)\big)=-\,c(j;x)=c(j;x).$
Step D: use (SR2) to pass from $j$ and $t^{2}$ to any involution. Let $a$ be any involution. By (SR2) applied to the involutions $(j,a)$ there is $t$ with $t^{2}=ja$, hence left–multiplying by $j$ (using $j^{2}=e$) gives $a=jt^{2}$. Therefore, for every $x$,
$c(j;xa)=c(j;xjt^{2})\stackrel{{\scriptstyle(ii)}}{{=}}c(j;xj)\stackrel{{\scriptstyle(i)}}{{=}}c(j;x).$
As $G$ is generated by involutions, this shows $c(j;x)$ is independent of $x$. Evaluating at $x=e$ gives $c(j;e)=f(j)$; hence
$f(xj)-f(x)=c(j;x)=f(j)\quad\text{for all }x\in G.$
That is, $f(xj)=f(x)+f(j)$, completing the case $q=1$.
Inductive step for $q\geq 2$. We now prove the additivity $f(xy)=f(x)+f(y)$ by induction on $q=\ell(y)$. The base case $q=1$ (i.e., $f(xj)=f(x)+f(j)$ for $j\in\mathcal{I}$) was completed above.
Assume the statement holds for all elements $Y\in G$ with $\ell(Y)=q-1$. That is, we assume $f(xY)=f(x)+f(Y)$ for any $x\in G$. Let $y\in G$ be an element with $\ell(y)=q\geq 2$. We can write $y$ as $y=Yj$, where $j\in\mathcal{I}$ and $\ell(Y)=q-1$. We then compute as follows:
$f(xy)$ $=f(x(Yj))=f((xY)j)$
$=f(xY)+f(j)$ (by applying the base case $q=1$ to $X=xY$)
$=\big(f(x)+f(Y)\big)+f(j)$ (by the induction hypothesis on $Y$)
$=f(x)+\big(f(Y)+f(j)\big)$ (since $H$ is abelian)
$=f(x)+f(Yj)$ (by applying the base case $q=1$ again, with $x=Y$)
$=f(x)+f(y)$
This completes the inductive step. By the principle of induction, the additivity $f(xy)=f(x)+f(y)$ holds for all $x,y\in G$.
Finally, we verify that $f$, now proven to be a homomorphism, also satisfies (J2). This follows from the homomorphism property and Lemma 2.3(I) (which itself was derived from (J1)):
$f(xy)+f(x^{-1}y)=\big(f(x)+f(y)\big)+\big(f(x^{-1})+f(y)\big)=\big(f(x)+f(x^{-1})\big)+2f(y)=2f(y).$
Thus, $f$ is a homomorphism that satisfies both Jensen equations. The proof of the theorem is complete. ∎
Corollary 2.9 (Parity normal form under (SR2)).
Assume $G$ satisfies (SR2) and let $f\in S_{1}(G,H)$ with $f(e)=0$. Then $f$ is a group homomorphism and:
$\exists\,u\in H[2]\ \text{such that}\quad f(i)=u\ \ \text{for every involution }i,$
and for any involution word $g=i_{1}\cdots i_{r}$,
$f(g)=\begin{cases}0,&r\ \text{even},\\ u,&r\ \text{odd}.\end{cases}$
Here, $H[2]:=\{h\in H:\ 2h=0\}$ denotes the $2$-torsion subgroup of $H$.
Proof.
Let $a,b$ be involutions. By (SR2), pick $t\in G$ with $t^{2}=ab$. Since $G$ is generated by involutions, $t$ is a product of involutions; hence, by Lemma 2.6, $2f(t)=0$. Using Lemma 2.3(I), we obtain
$f(ab)\;=\;f(t^{2})\;=\;2f(t)\;=\;0.$
Since $f$ is a homomorphism, $0=f(ab)=f(a)+f(b)$, hence f(b)=f(a)=f(a)f(b)=-f(a)=f(a) because $2f(a)=0$ (Lemma 2.3(I)). Thus all involutions have a common value $u\in H[2]$.
For any involution word $g=i_{1}\cdots i_{r}$, by additivity from the proof of Theorem 2.8, $f(g)=\sum_{k=1}^{r}f(i_{k})=r\cdot u,$ which equals 0 if $r$ is even and $u$ if $r$ is odd. The corollary follows. ∎
3. Applications
In this section, we demonstrate how the $(\mathrm{SR}_{2})$ criterion applies to several important classes of groups such as the symmetric and dihedral groups. We note that the results for these groups are already known (see [4]). Our contribution here is to show that these results can be recovered in a unified and direct manner as immediate consequences of our general $(\mathrm{SR}_{2})$ criterion. This highlights the structural insight provided by our approach.
3.1. Symmetric groups
We now show how our general criterion recovers the known result for the symmetric group. The fact that all Jensen solutions on $S_{n}$ are homomorphisms, and therefore have the parity form stated below, was first established by C.T. Ng [4] using properties of the quotient group $G/(\text{squares})$. A direct, combinatorial proof was later given by Trinh and Hieu [7]. Our contribution is to demonstrate that this result also follows as an immediate consequence of the $(\mathrm{SR}_{2})$ property.
Theorem 3.1 (Symmetric group).
Let $n\geq 2$ and let $\mathcal{I}$ be the set of transpositions in $S_{n}$. Then $(\mathrm{SR}_{2})$ holds for $S_{n}$ with respect to $\mathcal{I}$. Consequently, by Theorem 2.8,
$S_{1}(S_{n},H)=S_{1,2}(S_{n},H)\;\cong\;\mathrm{Hom}(S_{n},H).$
Moreover, every solution $f\in S_{1}(S_{n},H)$ with $f(e)=0$ is determined by a choice of $u\in H[2]$ and satisfies
$f(\sigma)=\begin{cases}0,&\text{if $\sigma$ even},\\ u,&\text{if $\sigma$ odd.}\end{cases}$
Proof.
Let $\tau_{1},\tau_{2}$ be transpositions in $S_{n}$ with $n\geq 2$. We show that there exists $t\in S_{n}$ with $t^{2}=\tau_{1}\tau_{2}$. There are three possibilities for the pair $(\tau_{1},\tau_{2})$.
Case 1: $\tau_{1}=\tau_{2}$. In this case, their product is $\tau_{1}\tau_{2}=\tau_{1}^{2}=e$, the identity element. The condition (SR2) requires finding $t\in S_{n}$ such that $t^{2}=e$. This is trivially satisfied by choosing $t=e$.
Case 2: $\tau_{1},\tau_{2}$ are distinct and intersect in a single element. Write $\tau_{1}=(ab)$ and $\tau_{2}=(bc)$ with $a,b,c$ pairwise distinct. Then
$\tau_{1}\tau_{2}=(ab)(bc)=(abc),$
a $3$-cycle. Let $t:=(acb)$; since the square of a $3$-cycle is its inverse,
$t^{2}=(acb)^{2}=(abc)=\tau_{1}\tau_{2}.$
Case 3: $\tau_{1},\tau_{2}$ are disjoint. Write $\tau_{1}=(ab)$ and $\tau_{2}=(cd)$ with $a,b,c,d$ pairwise distinct. Then
$\tau_{1}\tau_{2}=(ab)(cd).$
Let $t:=(acbd)$ be the $4$-cycle $a\mapsto c\mapsto b\mapsto d\mapsto a$. The square of a $4$-cycle is the product of the two opposite transpositions, hence
$t^{2}=(acbd)^{2}=(ab)(cd)=\tau_{1}\tau_{2}.$
In all three cases there exists $t\in S_{n}$ with $t^{2}=\tau_{1}\tau_{2}$, so $(\mathrm{SR}_{2})$ holds for $S_{n}$ with respect to the set of transpositions. By Theorem 2.8 we obtain $S_{1,2}(S_{n},H)\cong\mathrm{Hom}(S_{n},H)$. Finally, since $(S_{n})_{\mathrm{ab}}\cong C_{2}$ (the commutator subgroup is $A_{n}$) and $S_{n}$ is generated by involutions, Corollary 2.9 gives the stated parity formula with parameter $u\in H[2]$. ∎
3.2. Dihedral groups
Our criterion ($\mathrm{SR}_{2}$) also provides a particularly insightful perspective on the dihedral groups. While the general result that all Jensen solutions on $D_{m}$ are homomorphisms is known [4], the $(\mathrm{SR}_{2})$ property itself reveals a sharp dichotomy: it holds if and only if $m$ is odd. This provides a new structural explanation for why the odd-order case is particularly well-behaved under our approach.
Theorem 3.2 (Dihedral dichotomy).
Let $D_{m}=\langle r,s\mid r^{m}=e,\ s^{2}=e,\ srs=r^{-1}\rangle$ be the dihedral group, and let $\mathcal{I}=\{sr^{k}\mid 0\leq k<m\}$ be its set of reflections.
  1. (I)
    If $m$ is odd, then $D_{m}$ satisfies $(\mathrm{SR}_{2})$ with respect to $\mathcal{I}$. Consequently, $S_{1}(D_{m},H)=S_{1,2}(D_{m},H)=\mathrm{Hom}(D_{m},H).$
  2. (II)
    If $m$ is even, $(\mathrm{SR}_{2})$ may fail in general.
Proof.
(I) For reflections $sr^{i},sr^{j}$ we have $(sr^{i})(sr^{j})=r^{\,j-i}$. When $m$ is odd, $2$ is invertible mod $m$: choose $u$ with $2u\equiv 1\pmod{m}$. Then for every $k$ we have $r^{k}=(r^{ku})^{2}$, so products of two reflections are squares. Thus $(\mathrm{SR}_{2})$ holds and Theorem 2.8 applies.
(II) For $m$ even, not every rotation is a square. For example in $D_{4}$, $s\cdot(sr)=s^{2}r=r$, while the set of squares is $\{e,r^{2}\}$ (indeed, $(r^{a})^{2}=r^{2a}$ and $(sr^{a})^{2}=e$). Hence $r$ is not a square and $(\mathrm{SR}_{2})$ fails for the pair $(s,sr)$. ∎
Remark 3.3.
Theorem 3.2(II) established that the (SR2) criterion fails when $m$ is even, as not all rotations are squares. This raises the question of whether the conclusion of Theorem 2.8 (namely, $S_{1}(G,H)=\mathrm{Hom}(G,H)$) also fails. The following example confirms this by constructing an explicit solution to (J1) that is not a group homomorphism.
Example 3.4 (A non-homomorphic Jensen solution on $D_{2k}$ when $m=2k$ is even).
Let $D_{m}$ be as in Theorem 3.2 with $m=2k$ even, and let $(H,+)$ be an abelian group. Fix $u,c\in H$ with $2u=0$ and $2c=0$ (i.e. $u,c\in H[2]$). Define $f:D_{m}\to H$ by
$f(r^{2t})=0,\qquad f(r^{2t+1})=u,\qquad f(sr^{j})=c\quad(t\in\mathbb{Z},\ j\in\mathbb{Z}/m\mathbb{Z}).$
Then $f$ satisfies the Jensen equation (J1) for all $x,y\in D_{m}$, but $f$ is not a homomorphism provided $u\neq 0$.
Proof.
Well-definedness. Because $m$ is even, the parity of the exponent $j$ is well-defined modulo $m$; hence $f(r^{j})$ is unambiguous. Also $f(e)=f(r^{0})=0$.
Verification of (J1). Since $f$ takes values in $H[2]$, we have $2f(x)=0$ for all $x$. Thus (J1) is equivalent to
$f(xy)=f(xy^{-1})\qquad(\forall x,y\in D_{m}).$
We check this by cases, writing $y=r^{j}$ or $y=sr^{j}$ and $x=r^{p}$ or $x=sr^{p}$.
  • (i)
    $x=r^{p}$, $y=r^{j}$. Then $xy=r^{p+j}$ and $xy^{-1}=r^{p-j}$. Since $p+j\equiv p-j\pmod{2}$, both exponents have the same parity; hence $f(xy)=f(xy^{-1})$.
  • (ii)
    $x=r^{p}$, $y=sr^{j}$. Then $xy=r^{p}sr^{j}$ and $xy^{-1}=r^{p}r^{-j}s=r^{p-j}s$. Both products are reflections, so $f(xy)=f(xy^{-1})=c$.
  • (iii)
    $x=sr^{p}$, $y=r^{j}$. Then $xy=sr^{p+j}$ and $xy^{-1}=sr^{p-j}$, again both reflections; hence $f(xy)=f(xy^{-1})=c$.
  • (iv)
    $x=sr^{p}$, $y=sr^{j}$. Using $sr^{a}s=r^{-a}$, we get
    $xy=sr^{p}sr^{j}=r^{-p}r^{j}=r^{j-p},\qquad xy^{-1}=sr^{p}(r^{-j}s)=sr^{p-j}s=r^{-(p-j)}=r^{j-p}.$
    Thus $xy$ and $xy^{-1}$ are the same rotation (in particular have the same parity), and $f(xy)=f(xy^{-1})$.
In all cases $f(xy)=f(xy^{-1})$, so (J1) holds.
$f$ is not a homomorphism when $u\neq 0$. We have $f(r)=u$ and $f(s)=c$, while $sr$ is a reflection, so $f(sr)=c$. If $f$ were a homomorphism, we would need $f(sr)=f(s)+f(r)=c+u$, forcing $u=0$ (since $H$ is abelian). Hence for any $u\in H[2]\setminus\{0\}$ the map $f$ satisfies (J1) but is not additive. ∎
We conclude this section with two examples that generalize the preceding analysis and clarify the structure of the solution space identified in Theorem 2.8.
3.3. Consequences and examples
Example 3.5 (Coxeter-type situations).
Let $W$ be a Coxeter group generated by reflections. If for each rank–$2$ parabolic $W_{ij}\cong D_{m_{ij}}$ the order $m_{ij}$ is odd, then the argument of Theorem 3.2(1) applies in each rank–$2$ factor, so products of two reflections are locally squares and $(\mathrm{SR}_{2})$ can be verified by reduction to rank two. A full Coxeter-theoretic characterization of $(\mathrm{SR}_{2})$ is an interesting direction for future work.
Example 3.6 (Explicit solutions under $(\mathrm{SR}_{2})$).
Fix $H$ and choose any homomorphism $\phi:G_{\mathrm{ab}}\to H$. Then $f=\phi\circ\pi$ (with $\pi:G\to G_{\mathrm{ab}}$ the canonical projection) solves (J1) and (J2). Conversely, by Theorem 2.8 every solution arises this way.
4. Final remark: Comparison with complex-valued theory
and semigroup variants
When $H=\mathbb{C}$, Stetkær [6] proves
$f(xy)=f(x)+f(y)+\tfrac{1}{2}f([x,y])$
on $G/[G,[G,G]]$, and characterizes the odd-solution quotient by homomorphisms. Our criterion $(\mathrm{SR}_{2})$ can be seen as a group-structural hypothesis ensuring the commutator correction term vanishes without dividing by two in the codomain.
On semigroups with endomorphisms, recent results [2, 1] solve generalized Jensen equations assuming $H$ is $2$-torsion-free and at least one endomorphism is surjective; our results are orthogonal, as we require no endomorphisms but impose $(\mathrm{SR}_{2})$ on the source group.